OpenStudy (anonymous):

There are 9 items at the salad bar. How many ways can Gwen select 3 items from the salad bar? A. 3 B. 27 C. 84 D. 504 Question Resources

OpenStudy (anonymous):

@Arnab09 @ikatouni

OpenStudy (anonymous):

do you know combination?

OpenStudy (anonymous):

go through that topic carefully, you can easily understand then it's 9C3 = 84

OpenStudy (anonymous):

\[9-3=6~~~~~~~so,\]\[\color{red}{ \frac{9!}{3! \times 6!} }\] \[9!=6! \times 7 \times 8 \times 9\] and 6! cancels. YOU HAVE LEFT.\[\color{red}{ \frac{7 \times 8 \times 9}{1 \times 2 \times 3} }\]

OpenStudy (anonymous):

\[7 \times 4 \times 3=84\]